# Real-IAD Variety: Pushing Industrial Anomaly Detection Dataset to a Modern Era

### v1

@@ Abstract
Industrial Anomaly Detection (IAD) is critical for enhancing operational safety, ensuring product quality, and optimizing manufacturing efficiency across global industries. However, the IAD algorithms are severely constrained by the limitations of existing public benchmarks. Current datasets exhibit restricted category diversity and insufficient scale, frequently resulting in metric saturation and limited model transferability to real-world scenarios. To address this gap, we introduce Real-IAD Variety, the largest and most diverse IAD benchmark, comprising 198,960 high-resolution images across 160 distinct object categories. Its diversity is ensured through comprehensive coverage of 28 industries, 24 material types, and 22 color variations Our comprehensive experimental analysis validates the benchmark’s substantial challenge: state-of-the-art multi-class unsupervised anomaly detection methods experience significant performance degradation when scaled from 30 to 160 categories. Crucially, we demonstrate that vision-language models exhibit remarkable robustness to category scale-up, with minimal performance variation across different category counts, significantly enhancing generalization capabilities in diverse industrial contexts. The unprecedented scale and complexity of Real-IAD Variety position it as an essential resource for training and evaluating next-generation foundation models for anomaly detection. By providing this comprehensive benchmark with rigorous evaluation protocols across multi-class unsupervised, multi-view, and zero-/few-shot settings, we aim to accelerate research beyond domain-specific constraints, enabling the development of scalable, general-purpose anomaly detection systems. Real-IAD Variety will be made publicly available to facilitate innovation in this critical field.
@@ 1 Introduction
The field of anomaly detection has witnessed a paradigm shift towards fully unsupervised [1, 2, 3], multi-class unsupervised (unified) [4, 5, 6], and zero-/few-shot [7, 8, 9] learning frameworks, marking a transformative era in model design. Within the specialized domain of Industrial Anomaly Detection (IAD), substantial progress has been achieved in addressing the inherent complexity and variability of industrial processes.
@@
Figure 1: Industry distribution of the proposed Real-IAD Variety dataset. The dataset encompasses 8 major industrial categories (denoted as c): electrical, transport, cultural products, metal, general, electronics, rubber plastic, and other manufacturing sectors. These major categories are further subdivided into 28 industrial subcategories (denoted as s). Complete nomenclature details are provided in Appendix.
@@
Recent IAD research has predominantly concentrated on developing unified models capable of detecting anomalies across multiple heterogeneous industrial systems or processes [4, 10, 11, 12, 5]. In parallel, zero-shot and few-shot anomaly detection approaches [13, 7, 8, 9] have emerged to identify anomalies in previously unseen domains without requiring domain-specific prior knowledge. These advances are fundamentally transforming IAD, enabling the development of more robust, efficient, and adaptable detection frameworks.
@@
Despite the notable contributions of existing benchmarks such as MVTec AD [14], VisA [15], PAD [16], and Real-IAD [17], these datasets exhibit critical limitations in category diversity and scenario coverage, thereby constraining the comprehensive evaluation of contemporary unified and zero-/few-shot IAD models. Specifically, Real-IAD [17], which currently offers the largest category count, contains only 30 categories, while the widely adopted MVTec AD and VisA datasets comprise merely 15 and 12 categories, respectively. Although recent efforts [18, 19, 20] have attempted to synthesize large-scale anomaly images using pre-trained diffusion models, these approaches continue to face challenges in generating authentic and diverse anomalies.
@@
To advance IAD research toward unified and zero-/few-shot paradigms, there exists an urgent need for datasets that encompass substantially broader industrial categories and defect typologies. To address these limitations and propel IAD datasets into a new era, we introduce Real-IAD Variety, a large-scale benchmark that represents a significant advancement in the field. As illustrated in Figure 1 and Figure 3, this dataset substantially expands the scope of IAD data collection, featuring 160 categories spanning 28 industries, 24 material types, and 22 color variations, with a total of 198,960 images accompanied by pixel-level annotations. The category count of Real-IAD Variety is approximately 5.3 times that of Real-IAD [17], the previously largest IAD dataset. By dramatically increasing defect type diversity, Real-IAD Variety becomes the first multi-industry IAD dataset, thereby overcoming the inherent constraints of prior benchmarks for unified and zero-/few-shot evaluation.
Real-IAD Variety provides comprehensive coverage across 28 industries, including electrical manufacturing, transportation equipment, and cultural products, addressing the industry diversity gap present in Real-IAD, which primarily focuses on multi-view detection challenges. In contrast, existing datasets such as VisA [15] and MVTec AD [14] predominantly target narrower industrial sectors. Furthermore, Real-IAD Variety is designed to reflect real-world industrial requirements by incorporating 24 material types—including composite materials, metals, plastics, wood, and ceramics—offering substantially broader and more practical material representation compared to the limited ranges in Real-IAD [17], VisA [15], and MVTec AD [14].
The introduction of Real-IAD Variety not only augments training data diversity but also enables systematic investigation of the generalization capabilities of state-of-the-art (SOTA) IAD methods. We establish comprehensive benchmarks across three critical IAD settings: (1) multi-class unsupervised anomaly detection, (2) multi-view anomaly detection, and (3) zero-/few-shot anomaly detection. These settings are designed to reflect the diverse challenges encountered in real-world IAD applications, providing a rigorous framework for performance evaluation. Through extensive experimentation, we identify two key findings:
First, a substantial increase in category count leads to significant performance degradation in multi-class unsupervised anomaly detection (MUAD) methods, revealing that current MUAD algorithms exhibit limited scalability with respect to category diversity. For instance, as demonstrated in Table 3, when the category count increases from 30 to 60, 100, and finally 160, existing algorithms experience a performance decline of 1030%.
Second, zero-/few-shot learning approaches demonstrate greater robustness to category count variations, and when the number of categories exceeds a certain threshold, their performance surpasses that of MUAD methods. As shown in Table 4, pre-trained vision-language zero-/few-shot anomaly detection models maintain nearly consistent performance regardless of category count.
@@
  • We present Real-IAD Variety, a large-scale IAD dataset featuring unprecedented category diversity, extensive industry coverage, and high-quality pixel-level annotations. It expands the category count to 160 across 28 industries, 24 material types, and 22 color variations with 198,960 images, representing a substantial advancement over existing IAD benchmarks.
@@ 2.1 IAD Datasets
In Industrial Anomaly Detection (IAD), dataset diversity and scale are critical determinants of model effectiveness. Current 2D IAD benchmarks, including MVTec AD [14], VisA [15], MVTec Caption [21], and Real-IAD [17], have demonstrated considerable success in establishing foundational evaluation frameworks. Real-IAD [17], the largest existing dataset, encompasses 30 categories, while MVTec AD [14] and VisA [15] contain 15 and 12 categories, respectively. The MVTec Caption dataset [21] augments MVTec AD with descriptive textual annotations; however, it remains confined to a limited set of object types. Additional datasets such as MTD [22], MPDD [23], BTAD [24], and KolektorSDD [25] contribute valuable resources but are constrained by their limited scale and category diversity. Despite these contributions, existing datasets exhibit insufficient defect category variability and industrial domain coverage, thereby limiting their utility for evaluating unified and zero-/few-shot IAD paradigms. In contrast, our proposed Real-IAD Variety addresses these limitations by substantially enhancing category diversity, industrial scope, and defect typology, thereby advancing the field toward more comprehensive and realistic evaluation scenarios.
@@ 2.2 Unsupervised Anomaly Detection
Traditional unsupervised anomaly detection (UAD) algorithms have predominantly focused on single-class, single-view 2D images, training separate models for each object category using exclusively normal samples. These methods are typically categorized into three paradigms: (1) Discriminative methods (e.g., CutPaste [26], DRAEM [27], and SimpleNet [11]), which learn decision boundaries by distinguishing normal data from synthetically generated anomalies; (2) Reconstruction-based methods (e.g., DAE [28], O-CGAN [29], and RD [30]), which operate under the assumption that anomalous regions exhibit higher reconstruction errors compared to normal patterns; (3) Embedding-based methods (e.g., PatchCore [31], CFA [10], and CSFlow [32]), which model the distribution of normal feature representations in a frozen or learned embedding space. While these approaches have achieved notable success in single-class scenarios, their scalability to multi-class and cross-domain settings remains limited, motivating the development of unified detection frameworks.
@@ 2.3 Multi-class Unsupervised AD
To systematically evaluate the generalization capabilities of single-modality IAD models, the multi-class unsupervised anomaly detection (MUAD) setting was first introduced in UniAD [4]. This paradigm challenges models to train on all object categories within a dataset simultaneously, employing a single unified model for both training and inference. MUAD benchmarks encompass both traditional UAD methods, such as DRAEM [27], SimpleNet [11], CFA [10], CFLOW-AD [33], and RD [30]—and methods explicitly designed for unified training, including UniAD [4], OneNIP [5], MambaAD [12], DesTSeg [34], and Dinomaly [6]. However, existing MUAD evaluations have been conducted on datasets with limited category counts (typically $\leq$ 30 classes), leaving the scalability of these methods to large-scale, highly diverse industrial scenarios largely unexplored. Our Real-IAD Variety dataset, with its 160 categories, enables the first comprehensive assessment of MUAD performance under realistic industrial diversity.
@@ 2.4 Multi-view Anomaly Detection
2.5 Zero-/Few-Shot AD with VLMs
Figure 2: Data collection and annotation pipeline for the proposed Real-IAD Variety. The pipeline comprises a four-stage sequential process: (a) Material Preparation: This initial phase encompasses the assembly of a diverse array of materials, spanning 160 distinct categories sourced from 28 industrial domains and encompassing 24 material compositions. (b) Acquisition Equipment Design: The second phase involves the design of data capture apparatus, comprising one $\text{top}-\text{down}$ camera for overhead views and four peripheral cameras to capture lateral perspectives. (c) Data Collection and Annotation: The third phase pertains to the data collection process, which includes meticulous pixel-level manual annotation, rigorous algorithmic cross-validation, and iterative refinement. This process iterates until the model’s predicted Average Precision (AP) scores exhibit negligible variation below a predetermined threshold, following the methodology established in Real-IAD [17]. (d) Defect Taxonomy: The lower section illustrates 23 distinct defect types alongside their characteristic visual representations. Zoom in for enhanced visibility of defect regions delineated in red.
@@ 2.5 Zero-/Few-Shot AD with VLMs
The vision-language models (VLMs) have emerged as a promising direction for anomaly detection, enabling semantic reasoning and cross-domain generalization through natural language supervision. Recent advances have explored zero-shot (ZSAD) and few-shot (FSAD) anomaly detection using large-scale pre-trained models such as CLIP [36]. WinCLIP [37] pioneered the application of vision-language models for zero-shot anomaly classification and segmentation by introducing dual-class textual prompts and patch-wise window operations. AnomalyCLIP [7] employs object-agnostic textual prompts to capture generalizable normal and abnormal representations, further enhanced through auxiliary anomaly detection datasets. AdaCLIP [38] introduces learnable hybrid prompts and regional anomaly feature extraction to improve detection precision, while VCP-CLIP [39] designs Pre-VCP and Post-VCP modules to optimize textual embeddings with visual contextual prompts, promoting cross-modal information interaction.
@@
Recently, AdaptCLIP [8] treats CLIP as a foundational service, proposing alternating and comparative learning strategies based on three lightweight adapters to support zero- and few-shot generalization across domains. Despite these methodological advances, existing zero-/few-shot approaches have been validated primarily on small-scale datasets with limited category diversity (typically $\leq$ 30 classes), leaving their scalability and robustness under large-scale, multi-industry conditions unverified. Our Real-IAD Variety enables the first large-scale evaluation of vision-language anomaly detection methods across 160 categories, thereby facilitating systematic investigation of their generalization capabilities in realistic industrial settings.
3 Methodology: Real-IAD Variety
@@ 3.1 Data Collection Pipeline
Inspired by the methodology established in Real-IAD [17], we construct a rigorous three-stage data collection pipeline to ensure dataset comprehensiveness and annotation quality, as illustrated in Figure 2. The pipeline encompasses material preparation, acquisition equipment design, and data collection with iterative annotation refinement.
Stage 1: Diverse Material Preparation. To construct the Real-IAD Variety dataset, a team of 12 members dedicated 11,000 working hours to material selection and procurement, simulating a wide spectrum of real-world defects based on material characteristics. To enhance dataset representativeness, we assembled 160 object categories spanning 24 material types across 28 industrial domains. Representative samples are depicted in Figure 2a. Given that industrial production typically achieves yield rates exceeding 99%, naturally occurring defective parts are scarce. Therefore, leveraging extensive production line experience, we artificially introduced four defect variations for each material category, aggregating to 23 major defect types across the entire dataset.
@@
Stage 3: Data Collection and Annotation Refinement.
The data collection and annotation process for Real-IAD Variety adheres to the rigorous standards established by Real-IAD [17]. This stage involves meticulous pixel-level manual annotation, algorithmic cross-validation, and iterative refinement. The annotation process iterates until the model’s predicted Average Precision (AP) scores exhibit negligible variation below a predetermined threshold, ensuring annotation consistency and quality.

### v2

@@ Abstract
Industrial Anomaly Detection (IAD) is a cornerstone for ensuring operational safety, maintaining product quality, and optimizing manufacturing efficiency. However, the advancement of IAD algorithms is severely hindered by the limitations of existing public benchmarks. Current datasets often suffer from restricted category diversity and insufficient scale, leading to performance saturation and poor model transferability in complex, real-world scenarios. To bridge this gap, we introduce Real-IAD Variety, the largest and most diverse IAD benchmark. It comprises 198,950 high-resolution images across 160 distinct object categories. The dataset ensures unprecedented diversity by covering 28 industries, 24 material types, 22 color variations, and 27 defect types. Our extensive experimental analysis highlights the substantial challenges posed by this benchmark: state-of-the-art multi-class unsupervised anomaly detection methods suffer significant performance degradation (ranging from 10% to 20%) when scaled from 30 to 160 categories. Conversely, we demonstrate that zero-shot and few-shot IAD models exhibit remarkable robustness to category scale-up, maintaining consistent performance and significantly enhancing generalization across diverse industrial contexts. This unprecedented scale positions Real-IAD Variety as an essential resource for training and evaluating next-generation foundation IAD models.
@@ 1 Introduction
The field of anomaly detection has undergone a paradigm shift towards fully unsupervised [24, 37, 38], multi-class unsupervised (unified) [41, 11, 16], and zero-/few-shot [49, 10, 12] learning frameworks, ushering in a transformative era for model architecture. Within the specialized domain of Industrial Anomaly Detection (IAD), substantial progress has been achieved in addressing the inherent complexity and variability of industrial processes.
Figure 1: Visual illustration of diverse anomaly patterns in the Real-IAD Variety dataset. We present representative samples across 27 defect types, where anomaly regions are highlighted with red boundaries. Each sample is captured from five distinct viewpoints, demonstrating that certain defects are only detectable from specific angles while remaining invisible in others. Text annotations follow the format of “object (defect type)”, covering both structural damages (e.g., scratch) and logical inconsistencies (e.g., missing part).
@@
Figure 2: Industry distribution of the proposed Real-IAD Variety dataset. The dataset encompasses 8 major industrial groups (denoted as c): electrical, transport, cultural products, metal, general, electronics, rubber plastic, and other manufacturing sectors. These major categories are further subdivided into 28 industrial subcategories (denoted as s), and complete details are provided in Appendix.
@@
Recent IAD research has predominantly concentrated on developing unified models capable of detecting anomalies across multiple industrial objects using a single model [41, 25, 29, 17, 11]. Concurrently, zero-shot and few-shot anomaly detection approaches [13, 49, 10, 12] have emerged to identify anomalies in previously unseen domains without requiring task-specific fine-tuning. These advances are fundamentally reshaping IAD enabling more robust and adaptable frameworks. However, the progress is increasingly bottlenecked by the limitations of existing benchmarks.
@@
Despite the contributions of established benchmarks such as MVTec [1], VisA [50], PAD [48], and Real-IAD [39], these benchmarks exhibit critical limitations in category diversity and scenario coverage. For instance, Real-IAD [39], offers the largest category count, is limited to 30 categories, while the widely adopted MVTec and VisA datasets comprise only 15 and 12 categories, respectively. Therefore, it constrains the comprehensive evaluation of contemporary unified and zero-/few-shot IAD models. Furthermore, while synthetic data generation via pre-trained diffusion models has been explored [15, 35, 3], these approaches still face challenges in generating authentic and diverse anomalous patterns.
@@
To bridge this gap and propel IAD research toward unified and zero-/few-shot paradigms, there is an urgent requirement for benchmarks that encompass substantially broader industrial categories and defect typologies. In this paper, we introduce Real-IAD Variety, a large-scale benchmark that represents a significant advancement in the field. As illustrated in Figs. 1,  2 and 4, this dataset substantially expands the scope of IAD data collection, featuring 160 categories spanning 28 industries, 24 material types, 22 color variations and 27 defect types, culminating in a total of 198,950 images with high-fidelity pixel-level annotations. The category count of Real-IAD Variety is approximately 5.3 times that of its predecessor, Real-IAD [39], making it the first truly multi-industry benchmark that overcomes the constraints of prior datasets for unified and zero-/few-shot evaluation.
Beyond scale, Real-IAD Variety offers unprecedented diversity in materials (e.g., composites, metals, ceramics) and industries (e.g., electrical manufacturing, transport, electronics), addressing the practical needs of real-world deployment. We leverage this benchmark to establish rigorous evaluations across three critical settings: multi-class unsupervised anomaly detection (MUAD), multi-view anomaly detection, and zero-/few-shot anomaly detection.
Our extensive experiments yield two pivotal insights, scalability bottlenecks and paradigm robustness. We observe that an increase in category diversity leads to a significant performance degradation (10%-20%) in state-of-the-art MUAD methods, revealing their limited scalability when tasked with handling multiple category distributions. In contrast, zero-/few-shot approaches, particularly those leveraging large-scale vision-language models, demonstrate remarkable robustness. Their performance remains consistent regardless of category count, eventually surpassing MUAD methods once a certain complexity threshold is reached.
@@
  • We present Real-IAD Variety, a large-scale IAD dataset featuring unprecedented category diversity, extensive industry coverage, and high-quality pixel-level annotations. It expands the category count to 160 across 28 industries, 24 material types, 22 color variations and 27 defect types with 198,950 images, representing a substantial advancement over existing IAD benchmarks.
@@ 2.1 IAD Datasets
In Industrial Anomaly Detection (IAD), dataset diversity and scale are critical determinants of model effectiveness. Current 2D IAD benchmarks, including MVTec AD [1], VisA [50] and Real-IAD [39], have demonstrated considerable success in establishing foundational evaluation frameworks. However, their categorical scope remains limited. For example, Real-IAD [39], the largest existing dataset, encompasses 30 categories, while MVTec AD [1] and VisA [50] contain 15 and 12 categories, respectively. Additional datasets such as MTD [20], MPDD [22], BTAD [30], and KolektorSDD [36] contribute valuable resources but suffer from restricted scale and low categorical diversity.
Recent efforts have extended these benchmarks toward multi-modal or large-scale settings. For instance, MANTA [8] introduce descriptive textual annotations for multi-modal IAD focusing on tiny objects across 38 categories. In the logistics domain, Kaputt [19] presents a large-scale collection of 48,000 objects; however, its utility is constrained by the lack of pixel-level mask annotations, which are essential for precise defect localization. Furthermore, some recent benchmarks like COCO-AD [43] and ADNet [28] are constructed through the re-purposing or integration of existing datasets. COCO-AD [43] manually redefines categories from the COCO dataset as anomalies, which fundamentally differs from the authentic “defects” encountered in real-world industrial production. ADNet [28] primarily focuses on the consolidation of existing open-source benchmarks. Despite these contributions, existing datasets often exhibit limited category variability, insufficient scale, or a gap between synthetic anomalies and real-world industrial scenarios. In contrast, our proposed Real-IAD Variety addresses these limitations by substantially enhancing category diversity, industrial scope, and defect typology, thereby advancing the field toward more comprehensive and realistic evaluation scenarios.
@@ 2.2 Unsupervised Anomaly Detection
Traditional unsupervised anomaly detection (UAD) algorithms typically focuses on single-class, single-view 2D images, where models are trained independently for each object category using only normal samples. These methods are generally categorized into three paradigms: (1) Discriminative methods (e.g., CutPaste [26], DRAEM [42], and SimpleNet [29]), which learn decision boundaries by distinguishing normal data from synthetically generated anomalies; (2) Embedding-based methods (e.g., PatchCore [33], CFA [25], and CSFlow [34]), which model the distribution of normal feature representations in a frozen or learned embedding space. While these approaches have achieved notable success in single-class scenarios, their scalability to multi-class and cross-domain settings remains limited, motivating the development of unified detection frameworks; (3) Reconstruction-based methods (e.g., DAE [47], OCR-GAN [27], and RD [5]), which operate under the assumption that anomalous regions exhibit higher reconstruction errors compared to normal patterns.
@@ 2.3 Multi-Class Unsupervised AD
The Multi-class Unsupervised Anomaly Detection (MUAD) setting, pioneered by UniAD [41], requires a single unified model to handle training and inference across multiple categories simultaneously. MUAD benchmarks encompass both traditional UAD methods, such as DRAEM [42], SimpleNet [29], CFA [25], CFLOW-AD [14], and RD [5], explicitly designed for unified training, including UniAD [41], OneNIP [11], MambaAD [17], DesTSeg [45], LGC [9] and Dinomaly [16]. However, existing MUAD evaluations have been conducted on datasets with limited category counts (typically $\leq$ 30 classes), leaving the scalability of these methods to large-scale, highly diverse industrial scenarios largely unexplored. Our Real-IAD Variety dataset, with its 160 categories, provides the first rigorous testing ground for MUAD scalability.
@@ 2.4 Multi-View Anomaly Detection
2.5 Zero-/Few-Shot AD Models
@@
Vision-Language Models (VLMs) have emerged as a promising direction for anomaly detection, enabling semantic reasoning and cross-domain generalization. Recent advances have explored zero-shot (ZSAD) and few-shot (FSAD) anomaly detection using large-scale pre-trained models such as CLIP [32]. WinCLIP [21] introduces dual-class textual prompts and patch-wise window operation for ZSAD and FSAD. AnomalyCLIP [49] learns object-agnostic prompt embeddings to capture generalizable normal and abnormal representations through an auxiliary anomaly detection dataset. AdaCLIP [2] introduces learnable hybrid prompts and regional anomaly feature extraction to improve detection precision, while VCP-CLIP [31] designs Pre-VCP and Post-VCP modules to optimize textual embeddings with visual contextual prompts, promoting cross-modal information interaction. Recently, AdaptCLIP [10] treats CLIP as a foundational service, proposing alternating and comparative learning strategies based on three lightweight adapters to support zero- and few-shot generalization across domains.
@@
Unlike VLM-based ZSAD and FSAD, MetaUAS [12] is the first to reframe anomaly detection from a change detection perspective, achieving a pure-vision one-shot anomaly detection paradigm that eliminates the reliance on specific anomaly detection datasets for pre-training. Despite these methodological advances, existing ZSAD and FSAD approaches have been validated primarily on small-scale datasets with limited category diversity (typically $\leq$ 30 classes), leaving their scalability and robustness under large-scale, multi-industry conditions unverified. Our Real-IAD Variety dataset enables the first large-scale evaluation of ZSAD and FSAD methods across 160 categories, thereby facilitating systematic investigation of their generalization capabilities in realistic industrial settings.
Figure 3: Data collection and annotation pipeline for the proposed Real-IAD Variety. The pipeline comprises a four-stage sequential process: (a) Material Preparation: This initial phase encompasses the assembly of a diverse array of materials, spanning 160 distinct categories sourced from 28 industrial domains and encompassing 24 material compositions. (b) Acquisition Equipment Design: The second phase involves the design of data capture apparatus, comprising one top-down camera for overhead views and four peripheral cameras to capture lateral perspectives. (c) Data Collection and Annotation: The third phase pertains to the data collection process, which includes meticulous pixel-level manual annotation, rigorous algorithmic cross-validation, and iterative refinement. The process iterates until the variation in the model’s predicted Average Precision (AP) scores falls below a predetermined threshold, following the methodology established in Real-IAD [39]. (d) Defect Taxonomy: The lower section illustrates distinct defect types alongside their characteristic visual representations. Zoom in for enhanced visibility of defect regions delineated in red.
3 Real-IAD Variety
@@ 3.1 Data Collection Pipeline
Inspired by the methodology established in Real-IAD [39], we construct a rigorous three-stage data collection pipeline to ensure dataset comprehensiveness and annotation quality, as illustrated in Fig. 3. The pipeline encompasses material preparation, acquisition equipment design, and data collection with iterative annotation refinement.
Stage 1: Diverse Material Preparation. To construct the Real-IAD Variety dataset, a team of 12 members dedicated 11,000 working hours to material selection and procurement, simulating a wide spectrum of real-world defects based on material characteristics. To enhance dataset representativeness, we assembled 160 object categories spanning 24 material types across 28 industrial domains. Representative samples are depicted in Fig. 3a. Given that industrial production typically achieves yield rates exceeding 99%, naturally occurring defective parts are scarce. Therefore, leveraging extensive production line experience, we artificially introduced four defect variations for each material category, aggregating to 27 defect types across the entire dataset.
Table 1: Detailed Specifications of the Multi-view Acquisition System.
Specification Top-down Camera Peripheral Cameras
Sensor Resolution 5,328 $\times$ 3,040 4,096 $\times$ 3,000
Quantity 1 4
Mounting Angle 90 (Vertical) 40–45 (Oblique)
Lateral Accuracy 0.01 mm/pixel 0.028 mm/pixel
Working Distance 185 mm 310 mm
Field of View (FoV) 53 $\times$ 30.4 mm 114 $\times$ 84 mm
@@
Light Source RGBW Multi-spectral
[…]
Stage 3: Data Collection and Annotation Refinement. The data collection and annotation process for Real-IAD Variety adheres to the rigorous standards established by Real-IAD [39]. This stage involves meticulous pixel-level manual annotation, algorithmic cross-validation, and iterative refinement. The annotation process iterates until the model’s predicted Average Precision (AP) scores exhibit negligible variation below a predetermined threshold, ensuring annotation consistency and quality.